a) The resistor of an RC circuit has a resistance of 10 MΩ, the time constant is 1 s. What is the capacity of the capacitor? b) To what percent of the original value does the voltage of the capacitor decrease in 2 s?


1
Expert's answer
2021-06-28T16:55:01-0400

a)

"C=\frac{T}{2\pi R}=\frac{1}{2\cdot 3.14\cdot 10\cdot 10^6}=15.9~nF,"

b)

"R=\frac{T'}{2\pi}(\frac{1}{C'}-\frac 1C),\implies"

"C'=(\frac 1C+\frac{2\pi R}{T'} )^{-1}=6.36~nF,"

"p=1-\frac{C'}{C}=0.6=60\%."